OpenStudy (anonymous):

I have to find out why the serie 1/(n*(n+1)) is convergent. (further explanation inside)

OpenStudy (anonymous):

I have tried separating it like this: \[\frac{ 1 }{ n^2} + \frac{ 1 }{ n }\]

OpenStudy (anonymous):

1) compare to \(\frac{1}{n^2}\) will do it

OpenStudy (anonymous):

they are not equal

OpenStudy (anonymous):

you can separate it using partial fractions, which will actually give you the sum

OpenStudy (anonymous):

\[\frac{ 1 }{ n(n+1) } \neq \frac{ 1 }{ n^2 } + \frac{ 1 }{ n }\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes that is true, they are not equal

OpenStudy (anonymous):

ok, now i get what i did wrong. So the argument for saying that is convergent would be ?

OpenStudy (anonymous):

\[\frac{1}{n^2}+\frac{1}{n}=\frac{n+1}{n^2}\]

OpenStudy (anonymous):

\[\frac{1}{n(n+1)}=\frac{1}{n}-\frac{1}{n+1}\] is correct

OpenStudy (anonymous):

Yes, I understand why I cannot separate de denominator, but now I'm trying to understand why the serie is convergent

OpenStudy (anonymous):

reason that it converges is by comparison test you can say \(\frac{1}{n(n+1)}<\frac{1}{n^2}\) and as long as you know that \[\sum_{n=1}^\infty\frac{1}{n^2}\] is finite, then so is your sum

OpenStudy (anonymous):

hmmm, I understand now

OpenStudy (anonymous):

how would I calculate the sum ? Making an limit to the infinit ?

OpenStudy (anonymous):

infinite sorry

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty } \frac{ 1 }{ n(n+1) } = \frac{ 1 }{ \infty } = 0\] ?It converges at 0 ?